OpenStudy (anonymous):

Give an example of a sum of two rational expressions with different denominators, then perform the operation by showing all the steps, including how you found the common denominator. These rational expressions must have a variable in the denominator, such as (3x+1) / (x^2 – 1). Give another example of a sum, ratio, product, or different of two rational expressions for your classmates to solve. Again, the example must have a variable in the denominator. When might you use rational expressions in real life?

OpenStudy (anonymous):

3 x --- + ---- 2x-1 x+1 lcd = (2x+1)(x+1) now multiply each term by the lcd = 3(x+1) + x(2x-1) ----------------- (2x+1)(x+1) simplifying: = 3x + 3 + 2x^2 - x ---------------- (2x+1)(x+1) = 2x^2 + 2x + 3 ------------- (2x+1)(x+1)